every 20 minutes from 4:00 p.m. to 7:00 p.m., a commuter train crosses main street. for 3 minutes a gate stops cars from passing as the train goes by. what is the probability that a motorist approching the train crossing during this time interval will have to stop for the train?
Answer by Fombitz(32388)

Every 60 minutes, there is 9 minutes of potential stoppage time.
P=9/60=3/20